# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 6.5. The low $\kappa$ limit of $\lambda_{\rm max}$
where the default solution is understood, $x=\phi_{0}$, and the error $E_{z}=O(1+x^{\rho})e^{25|z|}$ is simply $O(1)e^{25z}$ since $x\in D$ (and $z>0$).
We now increase if need be the value of $c$ in $z=c\log\kappa^{-1}$ so that $|E_{z}|\leq\kappa^{-1/2}$. Using Proposition 5.4(4) to bound $f_{\rho}(x,z)$ above, we thus find that
$m_{\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor+1}\leq m_{-1,0}\cdot e^{-2z(1-\epsilon)}\big(1+O(1)\kappa^{1/2}\big)=m_{-1,0}\cdot\kappa^{2c(1-\epsilon)}\big(1+O(1)\kappa^{1/2}\big)\,$,
where $|O(1)|\leq 1$. From (71), we obtain
$m_{\lfloor z\kappa^{-1}\rfloor,\infty}\leq m_{-1,0}\,O(1)\kappa^{c_{0}-1}$ (72)
with $c_{0}=2c(1-\epsilon)$. Note that
$m_{-\infty,\lfloor z\kappa^{-1}\rfloor}\,=\,\sum_{i=\lfloor z\kappa^{-1}\rfloor}^{\infty}m_{-i-1,-i}$
@@
By Corollary 2.2, $n_{-i,-i-1}$ is equal to $m_{i,i+1}$ for the role-reversed ABMN solution $(b_{-i},a_{-i},n_{-i},m_{-i})$. Thus, from (72), we infer that $m_{-\infty,-\lfloor z\kappa^{-1}\rfloor}\leq m_{-1,0}\,O(1)\kappa^{c_{0}-1}$, whence also
@@
$m_{-\infty,\infty}\,=\,m_{-\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor}\Big(1-O(1)\kappa^{c_{0}}\Big)\,=\,m_{-1,0}\int_{-z}^{z}f_{\rho}(x,u)\,{\rm d}u\cdot\big(1+\kappa E_{z}\big)(1-O(1)\kappa^{c_{0}})$ (73)
where $E_{z}=O(1)e^{25|z|}$ (since $x^{\rho}=O(1)$, from $x\in D$). Given the selection of $c>0$ in the preceding proof, the choice $z=c\log\kappa^{-1}$ leads to $\kappa E_{z}=O(\kappa^{1/2})$.
@@
For $x\in(0,\infty)$, let $v\in\mathbb{R}$ be the value associated to $x$ by Proposition 5.4(2). Then
$\int_{\mathbb{R}\setminus[-z,z]}f_{\rho}(x,r)\,{\rm d}r\,=\,\frac{1}{f_{\rho}(1,-v)}\int_{\mathbb{R}\setminus[-z-v,z-v]}f_{\rho}(1,r)\,{\rm d}r\,.$
@@
$\int_{[-z,z]^{c}}f_{\rho}(x,r)\,{\rm d}r\,=\,\int_{[-z-v,z-v]^{c}}f_{\rho}(x,v+r)\,{\rm d}r\,=\,f_{\rho}(x,v)\int_{[-z-v,z-v]^{c}}f_{\rho}(1,r)\,{\rm d}r\,.$
Take r=vr=-v in Proposition 5.4(2,f) and use $f_{\rho}(x,0)=1$ (which is immediate from Definition 1.2) to find that $f_{\rho}(x,v)=1/f_{\rho}(1,-v)$. ∎
As $x$ varies over the precompact $D$, $v=v(x)\in\mathbb{R}$ remains bounded. So the factor $\frac{1}{f_{\rho}(-v)}$ is $O(1)$.
@@
with the constant $C$ absorbing the influence of the bounded offset $u$.
The integral $\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u$ is positive and finite, so
$\int_{-z}^{z}f_{\rho}(x,u)\,{\rm d}u\,=\,\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u\,\big(1-O(1)e^{-z}\big)\,$,
where we took $\epsilon\in(0,1/2)$. Since $z=c\log\kappa^{-1}$, we have $e^{-z}=\kappa^{c}$, so that (73) yields
$m_{-\infty,\infty}\,=\,\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u\cdot\big(1+O(1)\kappa^{1/2}\big)\big(1-O(1)\kappa^{c_{0}}\big)\big(1-O(1)\kappa^{c}\big)$
or simply $m_{-\infty,\infty}\,=\,m_{-1,0}\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u\,\cdot\big(1-O(1)\kappa^{c}\big)$ by decreasing the value of $c$ if need be.
@@
$n_{\infty,-\infty}\,=\,m_{-1,0}\int_{\mathbb{R}}g_{\rho}(x,u)\,{\rm d}u\,\cdot\big(1-O(1)\kappa^{c}\big)\,.$
@@
$\frac{n_{\infty,-\infty}}{m_{-\infty,\infty}}=\frac{\int_{\mathbb{R}}g_{\rho}(x,u)\,{\rm d}u\,\big(1-O(1)\kappa^{c}\big)}{\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u\,\big(1+O(1)\kappa^{c}\big)}\,.$
@@
$\frac{n_{\infty,-\infty}}{m_{-\infty,\infty}}=1+O(1)\kappa^{c}$
and we obtain (70) as desired. This completes the proof of Theorem 1.22. ∎
Remark. If we take $\rho\geq 1$, a more general error estimate (roughly $E_{z}=\exp\big\{(1+\rho)^{4}O(|z|)\big\}$) in Proposition 6.6 will lead to $c=c(\rho)\searrow 0$ as $\rho\to\infty$ in Theorem 1.22. The hypothesis $(\kappa,\rho)\in W$ is also needed, to enable $S_{\rho}$-tracking of the $s$-orbit, as in the remark that follows Proposition 6.3.
@@ 6.6. Scaled gameplay in the low-$\kappa$ limit
Here, the value $x=x_{\kappa}$ lies in the $(\kappa,\rho)$-central domain $D$ because the Nash equilibrium being played has battlefield zero. As such, $x_{\kappa}-1=O(\kappa)$ given the form of $D$ in Definition 1.20. Since $g_{\rho}(x,u)$ and $f_{\rho}(x,u)$ are smooth positive functions, $g_{\rho}(x_{\kappa},u)/f_{\rho}(x_{\kappa},u)=g_{\rho}(1,u)/f_{\rho}(1,u)\big(1+O(\kappa)\big)$.
@@
Proof of Theorem 1.18(2). Ethier and Kurtz’s [14, Corollary $7.4.2$] provides a framework for proving the convergence of discrete Markov chains to diffusion processes. For the framework to apply to a sequence of Markov chains $\{Y^{n}\}$ with transition kernels $p_{n}(x,\cdot)$, it is sufficient that the following conditions are met.
@@
The chains $Y^{n}$ may be specified on $0,\infty)$ rather than $\mathbb{N}$, by linear interpolation. When the above conditions are met, these chains are continuous real-valued processes on $0,\infty)$ whose scaled versions
$0,\infty)\to\mathbb{R}:u\to n^{-1}Y^{n}(n^{2}u)$
@@
We apply the framework with $Y^{n}(k)=n^{-1}X_{n^{-1},\rho}(n^{2}z,n^{2}k)$, so that $n\in\mathbb{N}$ corresponds to $\kappa$ in Theorem 1.18 via $n=\kappa^{-1}$. (It would seem that $\kappa$ must tend to zero through integer reciprocals. But in fact we may equally apply the framework with $n\to\infty$ in an arbitrary fashion.) To check that the framework is applicable, note that the scaled drift hypothesis is granted by Proposition 6.9 with $b(u)=\frac{1-S_{\rho}(1,u)^{\rho}}{1+S_{\rho}(1,u)^{\rho}}$. The magnitude of $Y_{n}$-jumps is $n^{-1}$, so $a_{n}(x)=1$ identically. By [43, Corollary $6.3.3$], the martingale problem for ${\rm d}X_{t}=a(X_{t}){\rm d}W_{t}+b(X_{t})\,{\rm d}t$ is well-posed when $a$ and $b$ are bounded with bounded continuous derivatives (in our case, $a=1$, and $b$ is smooth with $|b|\leq 1$).
@@
The outcome is the convergence asserted by Theorem 1.18(2), with the SDE-drift $R_{\rho}(u)$ given by $\frac{1-S_{\rho}(1,u)^{\rho}}{1+S_{\rho}(1,u)^{\rho}}$. The alternative formula claimed for $R_{\rho}(u)$ arises from the equality $S_{\rho}(1,u)^{\rho}=S_{1}(1,\rho^{2}u)$, which is precisely the identity noted after Lemma 5.1 with $x=1$. As is also noted there, $S_{1}(1,\rho^{2}u)\sim(8\rho^{2}|u|)^{{\bf 1}_{u<0}-{\bf 1}_{u>0}}$ as $|u|\to\infty$, which yields the asymptotics claimed for $R_{\rho}(u)$ when applied to the alternative formula. ∎
7. Directions
Our treatment has more or less directly posed certain open problems. These include formulating and solving $\rho$-Brownian Boost directly in continuous time via suitable classes of non-anticipatory strategies (see Section 1.6); proving negative results about solutions to ABMN when $(\kappa,\rho)$ lies high enough above the region $W$ specified in (8)—one could begin with $\kappa=1$ and $\rho>1$; and determining whether there are non-time-invariant equilibria in $\mathrm{TLP}(\kappa,\rho)$ and $\mathrm{BB}(\rho)$. [26, Section $7$] presents several directions for TKP($1,1$) including a discussion of the last problem. Here we indicate three broad directions for further study.
@@ 7. Directions
Finite-interval games. The Trail of Lost Pennies may be played on a finite interval $\llbracket-j,k\rrbracket$ for $j,k\in\mathbb{N}$. The game ends when $X$ reaches j-j or $k$ with terminal payments given by a quadruple $\big(m_{-j},m_{k},n_{-j},n_{k}\big)$. Under an analogue of the Nash-ABMN Theorem 1.8, time-invariant Nash equilibrium stake profiles would correspond to ABMN($\kappa,\rho$) elements that extend the boundary data to $\llbracket-j,k\rrbracket$. The finite-trail Mina margin map $\mathcal{M}_{\kappa,\rho}^{-j,k}:\phi_{0}\mapsto n_{k,-j}/m_{-j,k}$ satisfies the formula in Proposition 2.9(2) with summations over $\llbracket-j,k-1\rrbracket$ instead of $\mathbb{Z}$. The level sets of this map index equilibria of given Mina margin (or relative incentive) $n_{k,-j}/m_{-j,k}$. The finite-interval games were investigated for TLP($1,1$) in [26]. When the Mina margin is close to one, it appears that there is a unique equilibrium when $k-j\leq 5$; for kj=6k-j=6, there are three, and the number may be expected to grow as $2(k-j)+\Theta(1)$ for longer gameboards: see [26, Section $2.5$].
@@
We have not investigated the finite-interval games in this article, but the finite-trail Mina margin map offers a useful perspective on its results, with the low-$\rho$ convergence $\lambda_{\rm max}(\kappa,\rho)\to 1$ corresponding to $\mathcal{M}_{\kappa,\rho}^{-j,k}\to 1$ uniformly on compacts. The characteristic zigzag pattern seen in Figure 2.2 takes longer to appear as gameboard length rises when $\kappa$ is smaller: while $\mathcal{M}^{-9,9}_{1,1}=1$ has $27$ roots according to [26, Equation $(16)$], there are $21$ roots for $\mathcal{M}_{0.9,1}^{-9,9}=1$ as depicted in Figure 2.2. Likewise, the outset gameboard length for non-unique equilibria at given Mina margin may be expected rise as $\kappa$ drops: longer gameboards are needed at high-noise levels for the effects of stake turns to be felt.
@@
[…]
@@
Figure 7.1. A numerical approximation of the curve $\rho\to\lambda_{\rm max}(1,\rho)$ for $\rho$-values in $(0.8,1)$. The curve shown has been interpolated from a sequence of points $\big(\rho,\lambda_{\rm max}(\rho)\big)$, where each $\lambda_{\rm max}(\rho)$ is approximated by maximizing over a fine mesh the values in the central domain $D=D_{\kappa,\rho}$ of the finite-trail Mina margin map $\mathcal{M}_{j,k}$ for suitably high $j,k\in\mathbb{N}$.

### v4

@@ 6.5. The low $\kappa$ limit of $\lambda_{\rm max}$
where the default solution is understood, $x=\phi_{0}$, and the error $E_{z}=O(1+x^{\rho})e^{C_{1}|z|}$ is simply $O(1)e^{C_{1}z}$ since $x\in D$ (and $z>0$).
We now decrease if need be the value of $\eta$ in $z=\eta\log\kappa^{-1}$ so that $|E_{z}|\leq\kappa^{-1/2}$ (with $\eta<(2C_{1})^{-1}$). Using Proposition 5.4(4) to bound $f_{\rho}(x,z)$ above, we thus find that
$m_{\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor+1}\leq m_{-1,0}\cdot e^{-2z(1-\epsilon)}\big(1+O(1)\kappa^{1/2}\big)=m_{-1,0}\cdot\kappa^{2\eta(1-\epsilon)}\big(1+O(1)\kappa^{1/2}\big)\,$,
where $|O(1)|\leq 1$. From (6.20), we obtain
$m_{\lfloor z\kappa^{-1}\rfloor,\infty}\leq m_{-1,0}\,O(1)\kappa^{c_{0}-1}$ (6.21)
with $c_{0}=2\eta(1-\epsilon)$. Note that
$m_{-\infty,-\lfloor z\kappa^{-1}\rfloor}\,=\,\sum_{i=\lfloor z\kappa^{-1}\rfloor}^{\infty}m_{-i-1,-i}\,.$
@@
By Corollary 2.6, $n_{-i,-i-1}$ is equal to $m_{i,i+1}$ for the role-reversed ABMN solution $(b_{-i},a_{-i},n_{-i},m_{-i})$. Thus, from (6.21), we infer that $m_{-\infty,-\lfloor z\kappa^{-1}\rfloor}\leq m_{-1,0}\,O(1)\kappa^{c_{0}-1}$, whence also
@@
$m_{-\infty,\infty}\,=\,m_{-\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor}\Big(1-O(1)\kappa^{c_{0}}\Big)\,=\,m_{-1,0}\int_{-z}^{z}f_{\rho}(x,r)\,{\rm d}r\cdot\big(1+\kappa E_{z}\big)(1-O(1)\kappa^{c_{0}})$ (6.22)
where $E_{z}=O(1)e^{C_{1}|z|}$ (since $x^{\rho}=O(1)$, from $x\in D$). Given the selection of $\eta>0$ in the preceding proof, the choice $z=\eta\log\kappa^{-1}$ leads to $\kappa E_{z}=O(\kappa^{1/2})$.
@@
For $\xi\in(0,\infty)$, let $v\in\mathbb{R}$ be the value associated to $\xi$ by Proposition 5.4(2). Then
$\int_{\mathbb{R}\setminus[-z,z]}f_{\rho}(\xi,r)\,{\rm d}r\,=\,\frac{1}{f_{\rho}(1,-v)}\int_{\mathbb{R}\setminus[-z-v,z-v]}f_{\rho}(1,r)\,{\rm d}r\,.$
@@
$\int_{[-z,z]^{c}}f_{\rho}(\xi,r)\,{\rm d}r\,=\,\int_{[-z-v,z-v]^{c}}f_{\rho}(\xi,v+r)\,{\rm d}r\,=\,f_{\rho}(\xi,v)\int_{[-z-v,z-v]^{c}}f_{\rho}(1,r)\,{\rm d}r\,.$
Take u=vu=-v in Proposition 5.4(2,f) and use $f_{\rho}(\xi,0)=1$ (which is immediate from Definition 1.14) to find that $f_{\rho}(\xi,v)=1/f_{\rho}(1,-v)$. ∎
As $x$ varies over the precompact $D$, the quantity $v=v(x)\in\mathbb{R}$ remains bounded. So $\frac{1}{f_{\rho}(1,-v)}=O(1)$.
@@
with the constant $C$ absorbing the influence of the bounded offset $v$.
The integral $\int_{\mathbb{R}}f_{\rho}(x,r)\,{\rm d}r$ is positive and finite, so
$\int_{-z}^{z}f_{\rho}(x,r)\,{\rm d}r\,=\,\int_{\mathbb{R}}f_{\rho}(x,r)\,{\rm d}r\,\big(1-O(1)e^{-z}\big)\,$,
where we took $\epsilon\in(0,1/2)$. Since $z=\eta\log\kappa^{-1}$, we have $e^{-z}=\kappa^{\eta}$, so that (6.22) yields
$m_{-\infty,\infty}\,=\,m_{-1,0}\int_{\mathbb{R}}f_{\rho}(x,r)\,{\rm d}r\cdot\big(1+O(1)\kappa^{1/2}\big)\big(1-O(1)\kappa^{c_{0}}\big)\big(1-O(1)\kappa^{\eta}\big)$
or simply $m_{-\infty,\infty}\,=\,m_{-1,0}\int_{\mathbb{R}}f_{\rho}(x,r)\,{\rm d}r\,\cdot\big(1-O(1)\kappa^{\eta}\big)$ by decreasing the value of $\eta$ if need be.
@@
$n_{\infty,-\infty}\,=\,m_{-1,0}\int_{\mathbb{R}}g_{\rho}(x,r)\,{\rm d}r\,\cdot\big(1-O(1)\kappa^{\eta}\big)\,.$
@@
$\frac{n_{\infty,-\infty}}{m_{-\infty,\infty}}=\frac{\int_{\mathbb{R}}g_{\rho}(x,r)\,{\rm d}r\,\big(1-O(1)\kappa^{\eta}\big)}{\int_{\mathbb{R}}f_{\rho}(x,r)\,{\rm d}r\,\big(1+O(1)\kappa^{\eta}\big)}\,.$
@@
$\frac{n_{\infty,-\infty}}{m_{-\infty,\infty}}=1+O(1)\kappa^{\eta}$
and we obtain (6.19) as desired. This completes the proof of Theorem 1.23. ∎
Remark. If we take $\rho\geq 1$, a more general error estimate (roughly $E_{z}=\exp\big\{(1+\rho)^{4}O(|z|)\big\}$) in Proposition 6.6 will lead to $\eta=\eta(\rho)\searrow 0$ as $\rho\to\infty$ in Theorem 1.23. The hypothesis $(\kappa,\rho)\in W$ is also needed, to enable $S_{\rho}$-tracking of the $s$-orbit, as in the remark that follows Proposition 6.3.
@@ 6.6. Scaled gameplay in the low-$\kappa$ limit
The proposition is applied with the parameter $\xi$ set equal to the $\phi_{0}$-value (which we denote here by $x_{\kappa}$) associated to the given element $(a,b,m,n)\in$ ABMN($\kappa,\rho$). The result is applicable since $x_{\kappa}$ lies in the $(\kappa,\rho)$-central domain $D$ in view of the concerned Nash equilibrium having battlefield zero. And $x_{\kappa}\in D$ implies that $x_{\kappa}-1=O(\kappa)$ given the form of $D$ in Definition 1.21. Since $g_{\rho}(x,u)$ and $f_{\rho}(x,u)$ are smooth positive functions, $g_{\rho}(x_{\kappa},u)/f_{\rho}(x_{\kappa},u)=g_{\rho}(1,u)/f_{\rho}(1,u)\big(1+O(\kappa)\big)$.
@@
Proof of Theorem 1.18(2). Ethier and Kurtz’s [13, Corollary $7.4.2$] provides a framework for proving the convergence of discrete Markov chains to diffusion processes. For the framework to apply to a sequence of Markov chains $\{Y^{n}\}$ with transition kernels $p_{n}(x,\bullet)$, it is sufficient that the following conditions are met.
@@
The chains $Y^{n}$ may be specified on $0,\infty)$ rather than $\mathbb{N}$, by linear interpolation. When the above conditions are met, these chains are continuous real-valued processes on $0,\infty)$ whose speeded versions
$0,\infty)\to\mathbb{R}:u\to Y^{n}(n^{2}u)$
@@
Recall that, for $y\in\mathbb{R}$, $X_{\kappa,\rho}(y,\bullet):\mathbb{N}\to\mathbb{Z}$ denotes gameplay under the time-invariant Nash equilibrium of battlefield index zero, with initial condition $X_{\kappa,\rho}(0)=\lfloor y\rfloor$. We apply the framework with $Y^{n}(k)=n^{-1}X_{n^{-1},\rho}(nz,k)$ for $k\in\mathbb{N}$ (and with the domain being extended to $0,\infty)$ via interpolation). In this way, $n\in\mathbb{N}$ corresponds to $\kappa$ in Theorem 1.18 via $n=\kappa^{-1}$. (It would seem that $\kappa$ must tend to zero through integer reciprocals. But in fact we may equally apply the framework with $n\to\infty$ in an arbitrary fashion.) To check that the framework is applicable, note that the scaled drift hypothesis is granted by Proposition 6.9 with $b(u)=\frac{1-S_{\rho}(1,u)^{\rho}}{1+S_{\rho}(1,u)^{\rho}}$. Since the counter jump is $\pm 1$ at each step in $\mathrm{TLP}(\kappa,\rho)$, every jump of $Y^{n}$ has magnitude $n^{-1}$ and $a_{n}(x)=1$ identically. By [46, Corollary $6.3.3$], the martingale problem for ${\rm d}Z_{t}=a(Z_{t}){\rm d}W_{t}+b(Z_{t})\,{\rm d}t$ is well-posed when $a$ and $b$ are bounded with bounded continuous derivatives (in our case, $a=1$, and $b$ is smooth with $|b|\leq 1$).
@@
The outcome is the convergence asserted by Theorem 1.18(2), where in the SDE-drift $R_{\rho}(Z_{t})$, the function $R_{\rho}:\mathbb{R}\to(-1,1)$ is given by $R_{\rho}(u)=\frac{1-S_{\rho}(1,u)^{\rho}}{1+S_{\rho}(1,u)^{\rho}}$. The alternative formula claimed for $R_{\rho}(u)$ arises from the equality $S_{\rho}(1,u)^{\rho}=S_{1}(1,\rho^{2}u)$, which is precisely the identity noted after Lemma 5.1 with $\xi=1$. As is also noted there, $S_{1}(1,\rho^{2}u)\sim(8\rho^{2}|u|)^{{\bf 1}_{u<0}-{\bf 1}_{u>0}}$ as $|u|\to\infty$, which yields the asymptotics claimed for $R_{\rho}(u)$ when applied to the alternative formula. ∎
Epilogue: Directions
Our treatment has more or less directly posed certain open problems. We begin by surveying several such directions and then turn to some broader themes for further inquiry.
Invariance principles. For $\rho\in(0,1]$, we have exhibited Brownian Boost as a fine-mesh high-noise scaling limit of the Trail of Lost Pennies. How universal is this scaled structure? A natural direction is to seek to prove an invariance principle, a promising point of departure being given by variants of $\mathrm{TLP}(\kappa,\rho)$ whose stake and flip moves have symmetric jump distributions no longer supported on plus and minus one.
In another direction, our limit $\kappa\searrow 0$ represents one of high flip noise, with the contest rule fixed. Taking $\rho>0$ small in the contest success function $a^{\rho}/(a^{\rho}+b^{\rho})$ leads to outcomes close to those of a fair coin, just as making flip moves dominant does. Thus the limit $\rho\searrow 0$ for fixed $\kappa\in(0,1]$ yields a second high-noise regime whose scaling behaviour may be compared with the limit studied here. We have found a ‘Brownian Boost line’ along the positive $\rho$-axis; the proposed direction would seek a counterpart along the positive $\kappa$-axis.
A separate issue concerns the formulation of the model itself in continuous time. The chosen high-flip-noise representation of Brownian Boost has permitted us to exhibit its equilibrium structure explicitly, but it does so by circumventing the challenges inherent in instantaneous feedback loops. This may prompt attention to the problem of formulating and solving $\rho$-Brownian Boost directly via suitable classes of non-anticipatory strategies (see Section 1.4).
Discrete-game variants. Other problems concern how the solution theory we have developed breaks down in certain corners of parameter space. These include proving non-existence results for solutions of ABMN when $(\kappa,\rho)$ lies sufficiently above the region $W$ specified in (1.7)—one may begin with $\kappa=1$ and $\rho>1$—and determining whether non-time-invariant equilibria exist in $\mathrm{TLP}(\kappa,\rho)$ and $\mathrm{BB}(\rho)$. Several related directions for ${\rm TLP}(1,1)$ are discussed in [24, Section 7].
The ‘Hand It Over!’ one-step game discussed at the end of Chapter 2 may act as the single-step rule for a $\mathrm{TLP}(\kappa,\rho)$-type game played on integer intervals or on $\mathbb{Z}$. More generally, the Tullock contest and its ‘hand-over’ variant may be interpolated by introducing a pair of parameters in $[0,1]$ that specify the proportions of the winning and losing players’ stakes at a given turn that are transferred to the opponent, with the remainder surrendered to the bank. This yields a family of stake-governed tug-of-war games that interpolate between pure rent dissipation, at $(0,0)$, and pure rent transfer at $(1,1)$. It would be of interest to understand how equilibrium structure, incentive asymmetry, and discouragement effects vary across this family, and whether the high-noise scaling limits identified for $\mathrm{TLP}(\kappa,\rho)$ persist under such perturbations or undergo qualitative change.
Broader themes. The explicit ODE equilibria uncovered for Brownian Boost may also serve as a point of entry for analysts into questions of economic interest. Numerical investigations indicate unexpected intricacies in the quantified discouragement effect, while natural variants of Brownian Boost introduce higher-dimensional state variables and lead to multi-component coupled equations, posing challenges that may be of interest to PDE specialists.
We develop these broader themes through three directions for further study.
@@
Finite-interval games. The Trail of Lost Pennies may be played on a finite interval $\llbracket-j,k\rrbracket$ for $j,k\in\mathbb{N}$. The game ends when $X$ reaches j-j or $k$ with terminal payments given by a quadruple $\big(m_{-j},m_{k},n_{-j},n_{k}\big)$. Under an analogue of the Nash-ABMN Theorem 1.3, time-invariant Nash equilibrium stake profiles would correspond to ABMN($\kappa,\rho$) elements that extend the boundary data to $\llbracket-j,k\rrbracket$. The finite-trail Mina margin map $\mathcal{M}_{\kappa,\rho}^{-j,k}:\phi_{0}\mapsto n_{k,-j}/m_{-j,k}$ satisfies the formula in Proposition 2.9(2) with summations over $\llbracket-j,k-1\rrbracket$ instead of $\mathbb{Z}$. The level sets of this map index equilibria of given Mina margin (or relative incentive) $n_{k,-j}/m_{-j,k}$. The finite-interval games were investigated for TLP($1,1$) in [24]. When the Mina margin is close to one, it appears that there is a unique equilibrium when $k-j\leq 5$; for kj=6k-j=6, there are three, and the number may be expected to grow as $2(k-j)+\Theta(1)$ for longer gameboards: see [24, Section $2.5$].
@@
We have not investigated the finite-interval games in this article, but the finite-trail Mina margin map offers a useful perspective on its results, with the low-$\rho$ convergence $\lambda_{\rm max}(\kappa,\rho)\to 1$ corresponding to $\mathcal{M}_{\kappa,\rho}^{-j,k}\to 1$ uniformly on compacts. The characteristic zigzag pattern seen in Figure 2.2 takes longer to appear as gameboard length rises when $\kappa$ is smaller: while $\mathcal{M}^{-9,9}_{1,1}=1$ has $27$ roots according to [24, Equation $(16)$], there are $21$ roots for $\mathcal{M}_{0.9,1}^{-9,9}=1$ as depicted in Figure 2.2. Likewise, the outset gameboard length for non-unique equilibria at given Mina margin may be expected rise as $\kappa$ drops: longer gameboards are needed at high-flip-noise levels for the effects of stake turns to be felt.
@@
The map $(\kappa,\rho)\to\lambda_{\rm max}(\kappa,\rho)$. In (1.18), we extended the domain of $\lambda_{\rm max}$ by setting its values on the $\kappa=0$ axis equal to one. This accords with the absence of asymmetric equilibria in $\mathrm{BB}(\rho)$ due to Proposition 5.4(5). Regarding the second high-noise limit mooted above, where $\rho\searrow 0$ with $\kappa\in(0,1]$, it is reasonable to surmise analogously that $\kappa\mapsto\lambda_{\rm max}(\kappa,0)-1$ on $(0,1]$ also vanishes identically.
@@
Figure 6.1. A numerical approximation of the curve $\rho\to\lambda_{\rm max}(1,\rho)$ for $\rho$-values in $(0.8,1)$. The curve shown has been interpolated from a sequence of points $\big(\rho,\lambda_{\rm max}(\rho)\big)$, where each $\lambda_{\rm max}(\rho)$ is approximated by maximizing over a fine mesh the values in the central domain $D=D_{\kappa,\rho}$ of the finite-trail Mina margin map $\mathcal{M}_{1,\rho}^{j,k}$ for suitably high $j,k\in\mathbb{N}$.